OpenStudy (anonymous):

how do you get the (81)^3/4

OpenStudy (anonymous):

(81)^3/4 = \[\sqrt[4]{81}^{3}\]

OpenStudy (anonymous):

= 3^3 = 27

OpenStudy (anonymous):

gotcha